Opened 14 years ago

Closed 12 years ago

Last modified 6 years ago

#252 closed flight processing (fixed)

GB08/08, flight day 180b/2008, Wytham_Woods

Reported by: jatt Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_180b_Wytham_woods

Data arrived from ARSF via SATA N on 6th July 2009.

Scientific objective: see ~/arsf/arsf_data/in_progress/2008/ARSF_Applications/GB_and_NL/GB08-08 Science Case.pdf

Priority: α4H

PI: R Harding

Sensors:

  • Eagle (07/08/2009)
  • Hawk (07/08/2009)
  • Leica LIDAR (25/03/2010)
  • Digital Photography (05/01/2010)

Change History (25)

comment:1 Changed 14 years ago by jatt

Hawk files prefix renamed from Hawk to SWIRsh.

Hawk lines 6 and 8 are missing nav files.

For eagle and hawk lines 6 and 7 - files 6 and 7 are probably no good, looks as though they flown again as files 8 and 9.

comment:2 Changed 14 years ago by harg

  • Owner set to harg

comment:3 Changed 14 years ago by chrfi

  • eagle
    • line 1 sct -0.06 gpt -0.11
    • line 2 sct 0.29 gpt 0.18
    • line 3 sct -0.02 gpt -0.07
    • line 4 sct 0.09 gpt 0.00
    • line 5 sct 0.03 gpt -0.06
    • line 6 sct 0.07 gpt 0.06
  • hawk
    • line 1 sct 0.02 gpt 0.00
    • line 2 sct 0.00 gpt 0.00
    • line 3 sct 0.00 gpt 0.00
    • line 4 sct 0.10 gpt 0.06
    • line 5 sct -0.04 gpt -0.08
    • line 6 sct 0.08 gpt 0.04

comment:4 Changed 14 years ago by chrfi

  • Owner changed from harg to chrfi
  • Eagle
    • line 7 sct 0.78 gpt 0.08
  • Eawk
    • line 7 sct 0.62 gpt 0.02

comment:5 Changed 14 years ago by mark1

  • Owner changed from chrfi to mark1

Checking the delivery directory

comment:6 Changed 14 years ago by mark1

  • Owner changed from mark1 to chrfi

Delivery checked and made a few changes (cleaned up directory, edited readme file). Rsynced back to ~arsf/arsf_data/2009/flight_data/uk/GB08_08-2009_180b_Wytham_woods/

ready to deliver

comment:7 Changed 14 years ago by harg

This (with the 154a collection from the same project) has now (the afternoon of 07/08/2009) been delivered to Richard Harding from the Centre for Ecology and Hydrology. This included the raws as well as all the things listed at http://arsf-dan.nerc.ac.uk/trac/wiki/Procedures/EndUserDelivery.

comment:8 Changed 13 years ago by chrfi

  • Owner changed from chrfi to harg

comment:9 Changed 13 years ago by mark1

Photography delivered on hard disk 05/01/2010

comment:10 Changed 13 years ago by mark1

started lidar processing

comment:11 Changed 13 years ago by mark1

LIDAR data ready for delivery check at:
~airborne/workspace/GB08_08-2009_180b_Wytham_woods/delivery/20100322/GB08_08/

comment:12 Changed 13 years ago by emca

Delivery check complete.
Minor corrections made to read_me.
Changed PI on logsheet to R. Harding as per application form & ticket

Data is ready to deliver.

comment:13 Changed 13 years ago by mark1

  • Description modified (diff)

25/03/2010

comment:14 Changed 13 years ago by mark1

above comment refers to date of lidar delivery

comment:15 Changed 13 years ago by mark1

rsynced photography and LIDAR to main ~arsf repository.

comment:16 Changed 13 years ago by mark1

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

All data delivered and ready for archiving.

comment:17 Changed 12 years ago by knpa

  • Description modified (diff)
  • Owner changed from harg to knpa
  • Status changed from new to assigned

Started archiving.

comment:18 Changed 12 years ago by knpa

No webcam data.

No screenshots overlaid with vectors.

comment:19 Changed 12 years ago by knpa

renamed raw eagle and hawk files to standard.

comment:20 Changed 12 years ago by knpa

Created tarball, awaiting NEODC download.

comment:21 Changed 12 years ago by knpa

Backed up by NEODC.

Moved to archive disk.

comment:22 Changed 12 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:23 Changed 10 years ago by tipo

Nav data was showing in GraphNav as having the wrong date. Using DCALM.006 instead of DCALM.000 as the first file when extracting fixed it.

comment:24 Changed 6 years ago by asm

Archiving

Only reprocessed hyperspectral is still needed to be archived. Started.

comment:25 Changed 6 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.